You invest $1,000 at an annual rate of 6.75% compounded continuously.How long will it take for the value of the investment to double?
Answer by RAY100(1637)

formula for continuously compounding is ,,,A = Pe^rt
,
2000 = 1000 e^(.0675*t)
ln of both sides
ln2=.0675 *t,,,,,,,,,,,,,,,,,,,,,,,, lne^.0675t=.0675t
ln2 / .0675 = t
10.26 = t ,,,,,,,in yrs
,
check
2000 =1000 e^(.0675*10.26)=1000 * 2 = 2000,,,,,,ok